There is a Covid – 19 Pandemic in the world. This appeared in Ghana
and affected 5 people. Corvid – 19 affects 2 people a day. How long
does it take for there to be 1,000,000 covid – 19 affected people in
Ghana also how many covid – 19 people will there be on day 8
1
Expert's answer
2020-06-14T17:51:55-0400

"5+2n=1000000"

"n=\frac {1000000-5}{2}=499997.5"

Answer: "N=n+1=499998.5" days


"N_8=5+7\cdot2=19" people